What is the pH of a 0.16 M solution of Benzoic acid, pKa= 4.202
Answer #1

Initial concentration of benzoic acid = [C6H5COOH]initial = 0.16 M

pKa benzoic acid = 4.202

Ka = 10-pKa

Ka = 10-4.202

Ka = 6.28 x 10-5

ICE table C6H5COOH (aq) \rightleftharpoons C6H5COO- (aq) H+ (aq)
Initial conc. 0.16 M 0 0
Change -x +x +x
Equilibrium conc. 0.16 M - x +x +x

Ka = [C6H5COO-]eq[H+]eq / [C6H5COOH]eq

6.28 x 10-5 = [(x) * (x)] / (0.16 M - x)

Solving for x, x = 3.14 x 10-3 M

[H+] = x = 3.14 x 10-3 M

pH = -log[H+]

pH = -log(3.14 x 10-3 M)

pH = 2.50